Nicole Clothing manufactures embroidered jackets. It has reported the following standards for direct material and direct labor for the​ year:

Standard DM quantity per​ unit:            3 square feet

Standard cost per sq.​ foot:                                    ​$3.80/sq. ft

Standard DL quantity per​ unit:                2 hours per unit

Standard DL cost per​ hour:                                 ​$9.50/hr

Actual production was​ 14,000 jackets, with​ 33,600 total square feet used at an actual cost of​ $3.90/s                                               q. ft. Total direct labor hours incurred were​ 25,720 at a total cost of​ $249,484. Compute the direct material quantity variance.

A.

​$31,920 U

B.

​$32,760 F

C.

​$31,920 F

D.

​$32,760 U
Answer #1

Materials Quantity Variance

= (SQ-AQ) *SP

= (14,000*3 - 33,600) * 3.80

= (42,000 - 33,600) * 3.80

= 31,920 F